Status Report

ISS On-Orbit Status 8 Feb 2002

By SpaceRef Editor
February 8, 2002
Filed under , ,

All ISS systems continue to function nominally, except as noted
previously or below. Activities today emphasized science
research and auditing of supplies.

CDR Yuri Onufrienko started out by replacing the filters in two dust
collectors (PS1, PS2) in the FGB, without fan shutdown. This is a
recurring maintenance need.

Yuri also took care of the regular daily servicing duties, such as SM
life support systems maintenance, inspection of the BRPK-1 condensate
water separator and payload status check in the Lab module.

CDR downlinked the cardiovascular Velo-Ergometer measurements and gas
analyzer data which he had recorded on 12/28/01 on two floppy
diskettes as part of the Russian “Profilaktika” activity. The
transmission went to Moscow via S-band/OCA data file, by first
inserting the diskettes into the SSC (space station computer) VT disk
drive and then have it downloaded by MCC-Houston for subsequent
transfer to MCC-M.  Ground specialists stood by on the comm
loops during the activity.

Flight Engineer Carl Walz transferred accumulated exercise data from
the TVIS (treadmill) and RED (resistive exercise device) machines to
the MEC (medical equipment computer) for downloading to MCC-H at a
later time.

Flight Engineer Dan Bursch continued yesterdayâs activities on the
HRF (human research facility) Ultrasound equipment, today with a
physiological evaluation of the ultrasound life science gear,
including image downlink to the Huntsville POC (Payload Operations

Meanwhile, Carl performed a checkout of the new ZCG (Zeolite crystal
growth) furnace, which involved a brief activation of the ZCG
hardware to allow the ground an opportunity to verify data and
commanding prior to “upmassing” the actual science samples on Mission
8A.  ZCS operates in the ARIS-equipped EXPRESS Rack 2 (ER2).
 [Zeolites are substances that have a rigid crystalline
structure with a network of interconnected tunnels and cavities,
similar to a honeycomb. While a sponge needs to be squeezed in order
to release water, Zeolites give up their contents when they are
heated or under reduced pressure (the name "Zeolite" comes
from the Greek words zeo = to boil, and lithos = stone, literally
meaning "the rock that boils"). Zeolites have the ability
to absorb liquids and gases such as petroleum or hydrogen, while
remaining as hard as a rock. They form the backbone of the chemical
processes industry, and virtually all the world’s gasoline is
produced or upgraded using Zeolites. Industry wants to improve
Zeolite crystals so that more gasoline can be produced from a barrel
of oil, making the industry more efficient and thus reducing
America’s dependence on foreign oil].

Dan performed power-up and activation of the first of five growth
cylinders in the PCG-STES007 (protein crystal growth/single thermal
enclosure system unit 7), which is identical to the PCG-STES010
experiment already operating.  PCG-STES operates in ER4. [The
structural biology experiments conducted in the STES may provide a
basis for understanding the function of important macromolecules and
possibly contribute to the development of new biological
macromolecules. Their range includes proteins, polysaccharides and
other carbohydrates, lipids and nucleic acids of biological origin,
or those expressed in plant, animal, fungal or bacteria systems. The
fundamental goal for growing biological macromolecular crystals is to
determine their structure and the biological processes in which they
are involved. Macromolecules are selected, crystallized and analyzed
for atomic details — often by using X-ray crystallography.
Understanding these structures may impact the studies of medicine,
agriculture, the environment and other biosciences. After all, every
chemical reaction essential to life depends on the function of these

Crew conducted the carefully prepared life support
consumables/supply inventory audit.  Involved are such
categories as food, solid-waste containers (KTO), water containers
(EDV), towels and napkins (PHA), batteries, clothing, hygiene
equipment, photo items, etc. Ground teams have used crew reports via
notes, e-mails and calldowns to develop an up-to-date IMS (inventory
management system) database for each audit item.  Using the BCR
(bar code reader), items existing on board were recorded and matched
against tabular IMS data. Any improvements in manifesting the correct
amounts of consumables learned from the audit will fully benefit
plans for Expedition 6.  Resupply for Expedition 5 is already
packed for some categories, but transfer and trash plans are being
evaluated to make improvements.

FE-2 Bursch completed analysis of microbiological test samples taken
two days ago in the station with the SSK (surface sampler kit), WMK
(water microbiology kit) and MAS (microbial air sampler).  The
work involves visual analysis of covered media slides and Petri
dishes after their incubation period, including counting the number
of colonies visible and their estimated density. Results are recorded
first on data sheets and then entered in the MEC for eventual

Evaluation by MCC-H of the camcorder playback of yesterdayâs crew
inspection of the US Lab window showed numerous scratches and smudges
on the scratch pane (the innermost of the four double-layer glass
panes), observable with flashlight illumination against the night

Troubleshooting by MCC-M (TsUP) continues on one of the SNT-23
voltage and current stabilizers in the SM electrical power system,
which are designed to accept “external” power from the ARCUs
(American-to-Russian converter units) when an acute shortage arises
and a specific number of  storage batteries have already been
discharged.  It is not clear yet whether access through the TVIS
(treadmill) “pit” in the SM floor is required to do the
troubleshooting.  The short in the SNT-23 had shown itself by
tripping an RPCM (remote power controller module) in the Node.

After the resolution of the SNT problem, troubleshooting is also
required on the #2 ATU (audio terminal unit) in the Lab, which
tripped its RPCM last weekend.  The crew asked the ground for a
recovery date since there is only one additional ATU in the Lab.
 Also, a C&W (caution and warning) tone test activity
planned for this week was deferred until after the ATU #2 is

The VOA (volatile organics analyzer) was successfully commanded from
the ground to download ten data files contained in it.  To stop
the data file dump, a ground command to perform a calibration was
required. The data are being analyzed at present.

TsUP continues to investigate the integrity of certain waste
containers when exposed to vacuum.  Before the upcoming
separation of the Progress M1-256 (6P) from the SM aft port, Moscow
plans tS‡u¾pressurize its cargo compartment, and there was some
concern that liquid waste could be set free and affect the docking
mechanism.  If the analysis supports the concern, the waste
containers in question will be removed from the Progress and stowed.
RSC-Energia expects “positive results” of the tests.

TsUP is also dealing with a thermal system problem in the DC-1
docking compartment.  A heater appears to have “turned itself
on” and continues to operate. Compartment temperature is at 22 deg C,
and heater-off limit is several degrees below that value.

For the U.S. EVA from the Joint Airlock (JAL) on 2/20, the question
of final location of the Russian Strela adapter remains open.
Preliminary opinion at TsUP is that if Bursch and Walz are able to
separate the adapter from the U.S. grapple fixture after their
removal from the PMA-1 (pressurized mating adapter #1), it may be OK
to stow it at the FGB exterior.  If not, the combined assembly
would have to be taken inside.  A final resolution is expected
next Tuesday (2/12).

To troubleshoot the newly installed VTR2 (video tape recorder #2),
the ground plans to have the crew perform five VTR head cleaning
procedures next Monday through Friday, each day in the morning.
 If unsuccessful, Boeing will have a new VTR ready in time to be
launched on Mission 8A.

Update on Acoustic Noise:  Contrary to a previous report
(2/6), re-evaluation of the acoustic noise survey data taken by the
crew on 1/29 shows that the sound level situation onboard the Russian
segment (RS) has not degraded significantly from previous
expeditions. It appears that recordings from other areas were
incorrectly listed by Moscow as stemming from the crew quarters.
 However, the SM is still too noisy, i.e., its acoustic
environment is exceeding acceptable limits considerably even with the
modifications done to date. The same goes for the FGB and DC-1.
Clearly, continued work on the acoustics issue is necessary.

The Russian Uragan earth imaging program today focused on Sumatra
Island and Malay Peninsula (Kodak DCS, f/l 400 mm), Panama Canal
(DCS, 800 mm), and Mexico (LIV video cam).

CEO targets today were Industrialized Southeastern Africa
(after crossing the coast of South Africa, crew was to look well
to the left of track for oblique and limb shots of aerosol build-ups
over the Orange and Vaal River valleys)
, Angolan Biomass
(hot, dry weather now persists over southern Namibia
favorable for burning. Of interest: oblique views to the left or
right of track to document the number and spread of fires and smoke
plumes here)
, Rift Triple Junction, Ethiopia (other
than a few late day cumulus clouds, lighting was good to provide
shadows to enhance the relief associated with is complex of folds and
faults. Crew to look slightly to the right of track as ISS approached
the Red Sea from the SW)
, E. Mediterranean Dust and Smog
(with a strong storm system moving toward the eastern
Mediterranean this pass, crew to look to the left of track for plumes
of dust being swept northward from Egypt over the water)
, and
African Dust (satellite imagery continues to indicate dust
from NW Africa moving westward over the Atlantic. Crew was to take
advantage of the illumination oh:-is pass to look eastward to the
right of track for oblique and limb shots that document this
atmospheric phenomenon)

Upcoming Events:

  • 2/12: Readiness Review at JSC for ISS software transition to Release 2 (R2), a major step up.
  • 2/14: Transition to R2 to take place.
  • 2/15: Dry run and Readiness Review for EMU-EVA.
  • 2/20: EMU-EVA from JAL.
  • 2/21: Two reboost test maneuvers by Progress 6P (@ 1.5 m/s delta-V each).
  • TBD: Reboost by Progress 6P.
  • TBD: Separation of Progress 6P; ejection of Kolibri microsatellite.
  • 3/19-3/21: Launch of Progress M1-257 (7P)

U.S. and Russian Segment Status (as of 1:43 pm EST today):

Environmental Control and Life Support (ECLSS):

  • Elektron O2 generator is On (16 Amp mode). Vozdukh CO2 scrubber is in MANUAL cycle mode #5 (vacuum pump failed). ÊU.S. CDRA CO2 scrubber is Off.
  • BMP Harmful Impurities unit: Both absorbent beds (Filters #1 & #2) in Purify mode.
  • SM Working Compartment: Pressure (mmHg) — 748, temperature (deg C) — 25.8, ppO2 (mmHg) — 160.6, ppCO2 (mmHg) — 2.5 (?)
  • SM Transfer Compartment: Pressure (mmHg) — 743, temperature (deg C) — 20.8; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • FGB Cabin: Pressure (mmHg) — 744, temperature (deg C) — 21.3; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Node: Pressure (mmHg) — 738.16, temperature (deg C) — 22.7 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab: Pressure (mmHg) — 740.27, temperature (deg C) — 22.3, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a;
  • Joint Airlock: Pressure (mmHg) — Ê739.97, temperature (deg C) — 22.9.

(Partial pressures ppO2 and ppCO2 in U.S. segment (USOS) not available because MCA [major constituent analyzer] is failed and in Extended Life mode [= a state that preserves mass spectrometer vacuum but produces no pp data]).

Electrical Power Systems (EPS):

  • Beta Gimbal Assembly (BGA) 2B in Autotrack mode, BGA 4B in Directed position (125 degrees).
  • SM battery #2 is cycling; all other ( 7) SM batteries are in “partial charge” mode.
  • FGB battery #6 is cycling; battery #5 is offline; all other (4) batteries are in “partial charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Thermal Control Systems:

  • Air conditioner SKV-1 is Off (Freon leak). SKV-2 is On.

Command & Data Handling Systems:

  • C&C-2 MDM is prime, C&C-1 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is backup.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational.

Communications Systems:

  • S-band is operating nominally, and Ku-band is operating nominally in open loop pointing mode.
  • Audio subsystem operating nominally (except for ATU #2; see above).
  • Video subsystem operating nominally (but VTR2 is non-functional).
  • MCOR (medium-rate communications outage recorder) operating nominally.


  • SSRMS/Canadarm2 at Port stow position (on Keep Alive power on both strings).
  • RWS (robotics workstations) are Off.

ISS Orbit (as of this morning, 6:58 am EST):

  • Mean altitude — 386.2 km
  • Apogee — 390.2 km
  • Perigee — 382.2 km
  • Period — 92.3 min.
  • Inclination (to Equator) — 51.64 deg
  • Eccentricity — 0.0005913 (circle = 0.0)
  • Orbits per 24-hr. day — 15.60
  • Decay rate — 539 m (mean) in last 24 hours
  • Revolutions since FGB/Zarya launch (Nov. â98) — 18395
  • Solar Beta Angle — +4.2 deg (magnitude decreasing)
  • Current Flight Attitude — LVLH +XVV ZLV (local vertical/local horizontal: +X-axis in velocity vector; Z-axis in local vertical), with TEA (torque equilibrium attitude [pitch: -10 deg, yaw: -7 deg, roll: 0 deg]).

For more on ISS orbit and naked-eye visibility dates/times, see

SpaceRef staff editor.